# Relaxation-Free Min-$k$-Partition for PCI Assignment in 5G Networks

### v1

@@ I-A2 Optimization Methods
Beyond heuristics, several works formulate PCI assignment as a mathematical optimization problem and design algorithms to solve it. Since PCI assignment is inherently a difficult NP-hard combinatorial optimization problem, most algorithms encounter significant challenges in finding high-quality solutions. Consequently, simpler algorithms relying on greedy strategies and search methods are often employed to solve this problem, leading to suboptimal performance. Gui et al. [8] proposed a greedy algorithm to tackle the NP-hard binary quadratic programming (BQP) model for PCI assignment. Fairbrother et al. [16] modeled the PCI assignment problem as a two-level graph partitioning and presented a mixed-integer programming (MIP) model to solve it. Both studies incorporated weights to prioritize the minimization of mod-$3$ interference, mod-$30$ interference, collisions, and confusions.
While these methods account for mod-$3$ and mod-$30$ interference, they overlook the structural dependence of the objectives and the constraints, often struggling to balance solution quality and efficiency. This makes it difficult to scale them to large networks.
@@ III-C Graph Coloring for Quotient Assignment
The graph coloring method does not always guarantee that the PCI range constraint $30q_{i}+{\color[rgb]{0,0,0}r}_{i}\in\mathbb{Z}_{1008}$ is satisfied. When too many colors are used, some cells may be assigned overly large $q_{i}$, violating the PCI range constraint. To tackle this issue, we develop a heuristic post-processing procedure that re-assigns quotients $q_{i}$ to cells $i$ whose PCIs exceed $1007$, at the cost of potentially increasing the numbers of collisions and confusions. For each of these cells $i$, we fix its mod-$30$ value $r_{i}$, and update $q_{i}$ to a value $q\in\mathbb{Z}$ that satisfies $30q+r_{i}\leq 1007$ and minimizes the number of collisions and confusions with cells that already satisfy the PCI range constraint. We process the constraint-violating cells in descending order of the total number of neighbors and second-order neighbors, so as to prioritize highly constrained regions. The full algorithm is provided in Algorithm 1.
Result: Re-assigned quotient $\widetilde{\bm{q}}$.
for $r=0$ to $29$ do
       for $i\in\mathcal{I}_{r}$ do
      Sort $\mathcal{I}_{r}$ in decreasing order of $\{\deg_{v}\}_{v\in\mathcal{I}_{r}}$
       for $i\in\mathcal{J}_{r}$ do
             $\widetilde{q}_{i}\leftarrow q_{i}$
      for $i\in\mathcal{I}_{r}$ do
[…]
            [4pt] $\mathcal{J}_{r}\leftarrow\mathcal{J}_{r}\cup\{i\}$
return $\widetilde{\bm{q}}=\{\widetilde{q}_{i}\}_{i\in\mathcal{V}}$
@@ III-D Summary
Remark 1 (Comparison with existing works).
The works [25, 8] develop approaches for general mod-$k$ interference minimization for a single arbitrary $k$. Our framework can also be applied to that setting as follows. We first decompose a cell’s PCI value into $\text{PCI}_{i}=kq_{i}+r_{i}$, then solve the Min-$k$-Partition problem to determine the mod-$k$ value $r_{i}$ and minimize the mod-$k$ interference, and apply graph coloring to determine the quotient $q_{i}$ and eliminate collisions and confusions, leading to a valid PCI assignment.
Moreover, we emphasize that a major contribution of our work is a unified framework to tackle mod-$k$ interference for multiple $k$’s jointly. The main challenge in this setting is that the mod-$k$ interference for different $k$’s is correlated. For example, if two cells have mod-$3$ interference, then they must also have mod-$30$ interference. Our framework resolves this challenge through a CRT-based mod value decomposition. Furthermore, it extends beyond mod-$3$ and mod-$30$ values and applies to any mod value combinations that have the coprime structure required by the CRT.
@@ VI-B1 Experiment Setup
To evaluate algorithm performance under diverse topological conditions, we systematically vary both the network size and spatial density. Specifically, we consider network sizes $N\in\{100,200,500\}$, and connection radii $r\in\{0.1,0.2,0.3\}$ which correspond to sparse, medium, and dense graphs, respectively. For each configuration $\mathrm{RGG}(N,r)$, we generate 30 independent instances, which yields a total of 270 graph instances. For each type of graph, we compute standard structural metrics, including edge density, maximum clique size, average degree, and clustering coefficient. These statistics are summarized in Table I and allow us to interpret algorithm performance under varying structural complexities. The hyperparameters of GP-PMD is set as $\rho_{0}=10^{-8}$, $\gamma=1.1$, $\epsilon_{1}=10^{-5}$, and $\epsilon_{2}=10^{-10}$, which is the same configuration as [32]. We also conduct an ablation study to examine the effects of the hyperparameters on the performance of the algorithm, and the results are provided in Section A of the supplementary material.
@@ VI-B2 Experiment Results
We compare all baseline algorithms on the random geometric graphs using five evaluation metrics: mod-$3$ interference, mod-$30$ interference, number of collisions, number of confusions, and computational time. The overall performance is visualized in a heatmap (Fig. 4). The detailed box-plots are included in Section B of the supplementary material. The key observations are summarized as follows.
@@ Appendix D Proof of Theorem 2
Appendix A Ablation Study on Hyperparameters
@@
Appendix B Detailed Results on Synthetic Data
@@
Appendix C Detailed Results of statistical hypothesis tests

### v2

@@ I-A2 Optimization Methods
Beyond heuristics, several works formulate PCI assignment as a mathematical optimization problem and design algorithms to solve it. Since PCI assignment is inherently a difficult NP-hard combinatorial optimization problem, most algorithms encounter significant challenges in finding high-quality solutions. Consequently, simpler algorithms relying on greedy strategies and search methods are often employed to solve this problem, leading to suboptimal performance. Gui et al. [8] proposed a greedy algorithm to tackle the NP-hard binary quadratic programming (BQP) model for PCI assignment. Fairbrother et al. [16] modeled the PCI assignment problem as a two-level graph partitioning and presented a mixed-integer programming (MIP) model to solve it.
All existing optimization methods either incorporate weighted objectives to prioritize minimizing mod-$3$ interference, mod-$30$ interference, collisions, and confusions, or target a single arbitrary mod-$k$ interference component. As a result, they fail to account for the structural dependence between multiple modular interferences, which often leads to a trade-off between solution quality and computational efficiency. This limitation becomes especially problematic when scaling to large networks.
@@ III-C Graph Coloring for Quotient Assignment
The graph coloring method does not always guarantee that the PCI range constraint $30q_{i}+{\color[rgb]{0,0,0}r}_{i}\in\mathbb{Z}_{1008}$ is satisfied. When too many colors are used, some cells may be assigned overly large $q_{i}$, violating the PCI range constraint. To tackle this issue, we develop a heuristic post-processing procedure that re-assigns quotients $q_{i}$ to cells $i$ whose PCIs exceed $1007$, at the cost of potentially increasing the numbers of collisions and confusions. For each of these cells $i$, we fix its mod-$30$ value $r_{i}$, and update $q_{i}$ to a value $q\in\mathbb{Z}$ that satisfies $30q+r_{i}\leq 1007$ and minimizes the number of collisions and confusions with cells that already satisfy the PCI range constraint. We process the constraint-violating cells in descending order of the total number of neighbors and second-order neighbors, so as to prioritize highly constrained regions. The full algorithm is provided in Section A of the supplementary material [25].
@@ III-D Summary
While our main focus is the joint minimization of mod-$3$ and mod-$30$ interference, our decomposition framework naturally extends to arbitrary single mod-$k$ interference by expressing each PCI as $\text{PCI}_{i}=kq_{i}+r_{i}$, solving a Min-$k$-Partition for $r_{i}$, and applying graph coloring for $q_{i}$. Furthermore, it supports joint optimization over multiple moduli, provided they have the coprime structure required by the CRT. This yields a unified and scalable framework for systematically decomposing the multiple modular interferences minimization.
@@ VI-B1 Experiment Setup
To evaluate algorithm performance under diverse topological conditions, we systematically vary both the network size and spatial density. Specifically, we consider network sizes $N\in\{100,200,500\}$, and connection radii $r\in\{0.1,0.2,0.3\}$ which correspond to sparse, medium, and dense graphs, respectively. For each configuration $\mathrm{RGG}(N,r)$, we generate 30 independent instances, which yields a total of 270 graph instances. For each type of graph, we compute standard structural metrics, including edge density, maximum clique size, average degree, and clustering coefficient. These statistics are summarized in Table I and allow us to interpret algorithm performance under varying structural complexities. The hyperparameters of GP-PMD is set as $\rho_{0}=10^{-8}$, $\gamma=1.1$, $\epsilon_{1}=10^{-5}$, and $\epsilon_{2}=10^{-10}$, which is the same configuration as [32]. We also conduct an ablation study to examine the effects of the hyperparameters on the performance of the algorithm, and the results are provided in Section B of the supplementary material [25].
@@ VI-B2 Experiment Results
We compare all baseline algorithms on the random geometric graphs using five evaluation metrics: mod-$3$ interference, mod-$30$ interference, number of collisions, number of confusions, and computational time. The overall performance is visualized in a heatmap (Fig. 4). The detailed box-plots are included in Section C of the supplementary material [25]. The key observations are summarized as follows.
@@ Appendix D Proof of Theorem 2
Appendix A The full algorithm of the greedy quotient re-assignment for PCI range constraint satisfaction
Result: Re-assigned quotient $\widetilde{\bm{q}}$.
for $r=0$ to $29$ do
       for $i\in\mathcal{I}_{r}$ do
      Sort $\mathcal{I}_{r}$ in decreasing order of $\{\deg_{v}\}_{v\in\mathcal{I}_{r}}$
       for $i\in\mathcal{J}_{r}$ do
             $\widetilde{q}_{i}\leftarrow q_{i}$
      for $i\in\mathcal{I}_{r}$ do
[…]
            [4pt] $\mathcal{J}_{r}\leftarrow\mathcal{J}_{r}\cup\{i\}$
return $\widetilde{\bm{q}}=\{\widetilde{q}_{i}\}_{i\in\mathcal{V}}$
Appendix B Ablation Study on Hyperparameters
@@
Appendix C Detailed Results on Synthetic Data
@@
Appendix D Detailed Results of statistical hypothesis tests